Check my work 2 Coves Cakes is a local bakery. Price and cost Information follows: 1 13.71 points Price per cake Variable coRequired 1 Required 2 Required 3 Determine the bakerys margin of safety if it currently sells 420 cakes per month. (Round yoRequired 1 Required 2 Required 3 Determine the number of cakes that Cove must sell to generate $1,800 in profit. (Round your
Answer #1

1.

Selling price per unit = $13.71

Variable cost per unit = $3.53

Contribution margin per unit = Selling price per unit – Variable cost per unit

= 13.71 - 3.53

= $10.18

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 10.18/13.71

= 74.25%

Break even point (units) = Fixed cost/Contribution margin per unit

= 3,359.40/10.18

= 330 cakes

Break even sales in dollars = Fixed cost/Contribution margin ratio

= 3,359.40/74.25%

= $4,524

2.

Present sales = Number of units sold x Selling price per unit

= 420 x 13.71

= $5,758

Margin of safety in dollar sales = Present sales - Breakeven sales

= 5,758 - 4,524

= $1,234

3.

Units to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin per unit

= (3,359.40 + 1,800)/10.18

= 5,159.4/10.18

= 507 cakes